Find the square of the following numbers. 
86
Find the square of the following numbers. 
86

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(86)^2
=(90-4)^2
= (90)^2+ (4)^2- 2×90×4 [Since, (a+b)^2= a^2+b^2+2ab]​
= 8100+16–720
= 8116–720
= 7396

Thanks